# Distributionally Robust Beamforming and Estimation of Wireless Signals

### v1

@@ IV-B2 F-Norm Distributionally Robust Beamforming
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{x}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\operatorname{Tr}\left[{(\bm{R}_{x}-\hat{\bm{R}}_{x})^{\mathsf{H}}(\bm{R}_{x}-\hat{\bm{R}}_{x})}\right]\leq\epsilon^{2}_{1}\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},\end{array}$ (40)
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{x},\bm{V},\bm{U}}&\operatorname{Tr}\big{[}-\bm{V}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\left[\begin{array}[]{cc}\bm{V}&\bm{R}^{\mathsf{H}}_{xs}\\ \bm{R}_{xs}&\bm{R}_{x}\end{array}\right]\succeq\bm{0}\\ &\operatorname{Tr}\left[{\bm{U}}\right]\leq\epsilon^{2}_{1},\\ &\left[\begin{array}[]{cc}\bm{U}&(\bm{R}_{x}-\hat{\bm{R}}_{x})^{\mathsf{H}}\\ (\bm{R}_{x}-\hat{\bm{R}}_{x})&\bm{I}_{N}\end{array}\right]\succeq\bm{0},\\ &\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},~{}\bm{V}=\bm{V}^{\mathsf{H}},~{}\bm{U}=\bm{U}^{\mathsf{H}}\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{V}\succeq\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (41)
@@
As for the maximization over $\bm{R}_{xs}$ that cannot be characterized using the Wasserstein distance, we have
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{xs}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&\operatorname{Tr}\left[{(\bm{R}_{xs}-\hat{\bm{R}}_{xs})^{\mathsf{H}}(\bm{R}_{xs}-\hat{\bm{R}}_{xs})}\right]\leq\epsilon^{2}_{2},\end{array}$ (42)
which can be solved using, e.g., the projected gradient descent method, because it is a convex program. Note that the gradient of the objective function with respect to $\bm{R}_{xs}$ is $-2\bm{R}^{-1}_{x}\bm{R}_{xs}$.
@@ IV-C Multi-Frame Case: Dynamic Channel Evolution
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{Wiener-MF}}&=[\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}}][\bm{R}_{x}+\lambda\bm{I}_{N}]^{-1}\\ &=[\bm{P}\bm{H}^{\mathsf{H}}+\lambda\bm{W}^{\prime\mathsf{H}}][\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R}+\lambda\bm{I}_{N}]^{-1},\end{array}$ (44)
where $\lambda\geq 0$ is a tuning parameter to control the similarity between $\bm{W}$ and $\bm{W}^{\prime}$. Specifically, if $\lambda$ is large, $\bm{W}$ must be close to $\bm{W}^{\prime}$; if $\lambda$ is small, $\bm{W}$ can be far away from $\bm{W}^{\prime}$. normal-$\square$
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{x},\bm{R}_{xs},\bm{R}_{s}}&\operatorname{Tr}\big{[}-(\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}})^{\mathsf{H}}\cdot(\bm{R}_{x}+\lambda\bm{I}_{N})^{-1}\cdot\\ &\quad\quad\quad(\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}})+(\bm{R}_{s}+\lambda\bm{W}^{\prime}\bm{W}^{\prime\mathsf{H}})\big{]}\\ \text{s.t.}&d_{1}(\bm{R}_{x},{\hat{\bm{R}}}_{x})\leq\epsilon_{1},\\ &d_{2}(\bm{R}_{xs},{\hat{\bm{R}}}_{xs})\leq\epsilon_{2},\\ &d_{3}(\bm{R}_{s},{\hat{\bm{R}}}_{s})\leq\epsilon_{3},\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},\\ &\bm{R}_{s}\succeq\bm{0},~{}\bm{R}_{s}=\bm{R}^{\mathsf{H}}_{s},\end{array}$ (45)
whose objective function is monotonically increasing in $\bm{R}_{x}$ and $\bm{R}_{s}$.
The remaining distributional robustness modeling and analyses against the uncertainties in $\bm{R}_{x}$, $\bm{R}_{xs}$, $\bm{P}$, and $\bm{R}$ are technically straightforward, and therefore, we omit them here. Upon using the diagonal loading method on $\hat{\bm{R}}_{x}$, a distributionally robust beamformer for the multi-frame case is
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR-Wiener-MF}}=[\bm{P}\bm{H}^{\mathsf{H}}+\lambda\bm{W}^{\prime\mathsf{H}}]\cdot\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad[\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R}+\lambda\bm{I}_{N}+\epsilon_{1}\bm{I}_{N}]^{-1},\end{array}$
where $\epsilon_{1}$ is an uncertainty quantification parameter for $\hat{\bm{R}}_{x}$.
@@ V-A1 General Framework and Concrete Examples
$\min_{\bm{W}\in\mathbb{R}^{2M\times L}}\max_{\mathbb{P}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}\in\mathcal{U}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}}\operatorname{Tr}\mathbb{E}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}[\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}][\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}]^{\mathsf{T}},$ (46)
which is linear in $\bm{W}$. The proposition below gives the reformulation and solution of (46).
@@
Then, the distributionally robust $\underline{\bm{x}}$-nonlinear estimation problem (46) can be rewritten as a distributionally robust beamforming problem as
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{\underline{z}}\bm{W}^{\mathsf{T}}-\bm{W}\bm{R}_{\underline{zs}}-\bm{R}^{\mathsf{T}}_{\underline{zs}}\bm{W}^{\mathsf{T}}+\bm{R}_{\underline{s}}\big{]}\\ \text{s.t.}&d_{1}(\bm{R}_{\underline{z}},\hat{\bm{R}}_{\underline{z}})\leq\epsilon_{1},\\ &d_{2}(\bm{R}_{\underline{zs}},\hat{\bm{R}}_{\underline{zs}})\leq\epsilon_{2},\\ &d_{3}(\bm{R}_{\underline{s}},\hat{\bm{R}}_{\underline{s}})\leq\epsilon_{3},\\ &\bm{R}_{\underline{z}}\succeq\bm{0},~{}\bm{R}_{\underline{z}}=\bm{R}^{\mathsf{T}}_{\underline{z}},\\ &\bm{R}_{\underline{s}}\succeq\bm{0},~{}\bm{R}_{\underline{s}}=\bm{R}^{\mathsf{T}}_{\underline{s}},\end{array}$ (47)
where $\hat{\bm{R}}_{\underline{z}}\coloneqq\frac{1}{L}\bm{K}^{2}$, $\hat{\bm{R}}_{\underline{zs}}\coloneqq\frac{1}{L}\bm{K}\underline{\bm{S}}^{\mathsf{T}}$, $\hat{\bm{R}}_{\underline{s}}\coloneqq\frac{1}{L}\underline{\bm{S}}\underline{\bm{S}}^{\mathsf{T}}$, and $\underline{\bm{S}}\coloneqq[\operatorname{Re}\bm{S};~{}\operatorname{Im}\bm{S}]=[\underline{\bm{s}}_{1},\underline{\bm{s}}_{2},\ldots,\underline{\bm{s}}_{L}]$. In addition, the strong min-max property holds for (47): i.e., the order of $\min$ and $\max$ can be exchanged provided that the first three constraints are convex. As a result, given every pair of $(\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}})$, the optimal Wiener beamformer is [^4] For numerical stability in the inversion of $\bm{K}$, $L$ should not be too large.
[…]
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{T}}_{\underline{zs}}\bm{R}^{-1}_{\underline{z}}\bm{R}_{\underline{zs}}+\bm{R}_{\underline{s}}\big{]}\\ \text{s.t.}&d_{1}(\bm{R}_{\underline{z}},\hat{\bm{R}}_{\underline{z}})\leq\epsilon_{1},\\ &d_{2}(\bm{R}_{\underline{zs}},\hat{\bm{R}}_{\underline{zs}})\leq\epsilon_{2},\\ &d_{3}(\bm{R}_{\underline{s}},\hat{\bm{R}}_{\underline{s}})\leq\epsilon_{3},\\ &\bm{R}_{\underline{z}}\succeq\bm{0},~{}\bm{R}_{\underline{z}}=\bm{R}^{\mathsf{T}}_{\underline{z}},\\ &\bm{R}_{\underline{s}}\succeq\bm{0},~{}\bm{R}_{\underline{s}}=\bm{R}^{\mathsf{T}}_{\underline{s}}.\end{array}$ (49)
@@
The key is to treat $\bm{\varphi}(\underline{\mathbf{x}})$ as a new random vector $\underline{\mathbf{z}}\in\mathbb{R}^{L}$. For details, see Appendix G of the online supplementary materials. $\square$
In (47), $d_{1}$, $d_{2}$, and $d_{3}$ define three matrix similarity measures to quantify the uncertainties in $\underline{\hat{\bm{R}}_{z}}$, $\underline{\hat{\bm{R}}_{zs}}$, and $\underline{\hat{\bm{R}}_{s}}$, respectively; $\epsilon_{1}\geq 0$, $\epsilon_{2}\geq 0$, and $\epsilon_{3}\geq 0$ quantify the uncertainty levels. Proposition 3 reveals the benefit of the kernel trick (2), that is, the possibility to represent a nonlinear estimation problem as a linear one.
@@
which upper bounds the true estimation error covariance. normal-$\square$
@@
By using
$\hat{\bm{R}}_{\underline{z}}-\epsilon_{1}\bm{I}_{L}\preceq\bm{R}_{\underline{z}}\preceq\hat{\bm{R}}_{\underline{z}}+\epsilon_{1}\bm{I}_{L},$
@@
$\bm{\phi}^{\star}(\mathbf{x})=\bm{\Gamma}_{M}\cdot\frac{1}{L}\underline{\bm{S}}\bm{K}\cdot\left(\frac{1}{L}\bm{K}^{2}+\epsilon_{1}\bm{I}_{L}\right)^{-1}\cdot\bm{\varphi}(\underline{\mathbf{x}}),$ (52)
@@
$\begin{array}[]{l}\displaystyle\min_{\bm{W}\in\mathbb{R}^{2M\times L}}\operatorname{Tr}\mathbb{E}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}[\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}][\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}]^{\mathsf{T}}+\\ \quad\quad\quad\quad\quad\quad\quad\quad\epsilon_{1}\cdot\operatorname{Tr}[\bm{W}\bm{W}^{\mathsf{T}}],\end{array}$ (53)
which can be proven by replacing $\bm{R}_{\underline{z}}$ in (47) with its upper bound. normal-$\square$
@@
The kernelized eigenvalue thresholding method can be designed in analogy to Example 3. The two key steps are to obtain the eigenvalue decomposition of $\hat{\bm{R}}_{\underline{z}}=\bm{K}^{2}/L$ and then lift the eigenvalues; cf. (29). normal-$\square$
@@
${\hat{\bm{R}}}_{\underline{z}}-\epsilon\bm{I}_{L}\preceq{\bm{R}}_{\underline{z}}\preceq{\hat{\bm{R}}}_{\underline{z}}+\epsilon\bm{I}_{L},$
@@
${\hat{\bm{R}}}_{\underline{z}}-\epsilon\bm{F}\preceq{\bm{R}}_{\underline{z}}\preceq{\hat{\bm{R}}}_{\underline{z}}+\epsilon\bm{F}$
is employed for some $\bm{F}\succeq\bm{0}$.
@@
This is the first time to give the kernel ridge regression an interpretation of distributional robustness. The usual choice of $\bm{F}$ in Theorem 3 is the $L$-divided kernel matrix $\bm{K}/L$; see, e.g., [31, Eq. (4)], [24, Eqs. (15.110) and (15.113)]. As a result, from (52), we have
$\bm{\phi}^{\star}(\mathbf{x})=\bm{\Gamma}_{M}\cdot\underline{\bm{S}}\cdot\left(\bm{K}+\epsilon_{1}\bm{I}_{L}\right)^{-1}\cdot\bm{\varphi}(\underline{\mathbf{x}}),$ (54)
which is another type of kernel ridge regression (i.e., a new kernelized diagonal loading method).
@@
$\begin{array}[]{l}\displaystyle\min_{\bm{W}\in\mathbb{R}^{2M\times L}}\operatorname{Tr}\mathbb{E}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}[\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}][\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}]^{\mathsf{T}}+\\ \quad\quad\quad\quad\quad\quad\quad\quad\lambda\cdot\|\bm{W}\|^{2},\end{array}$ (55)
@@
Consider the nonlinear regression problem in Theorem 3. Its data-perturbed counterpart can be constructed by taking into account the data perturbation vectors $(\mathbf{\Delta}_{\underline{s}},\mathbf{\Delta}_{\underline{z}})$. Suppose that $\mathbf{\Delta}_{\underline{z}}$ is uncorrelated with $\underline{\mathbf{z}}$, with $\underline{\mathbf{s}}$, and with $\mathbf{\Delta}_{\underline{s}}$; in addition, $\mathbf{\Delta}_{\underline{s}}$ is uncorrelated with $\underline{\mathbf{z}}$. If the second-order moment of $\mathbf{\Delta}_{\underline{z}}$ is upper bounded by $\epsilon\bm{I}_{L}$, then the distributionally robust estimator of $\bm{W}$ becomes a kernel ridge regression (i.e., squared-$F$-norm regularized) method (53). The regularization term becomes $\operatorname{Tr}\big{[}\bm{W}\bm{F}\bm{W}^{\mathsf{H}}\big{]}$, which is known as the Tikhonov regularizer, if the second-order moment of $\mathbf{\Delta}_{\underline{z}}$ is upper bounded by $\epsilon\bm{F}$ for some $\bm{F}\succeq\bm{0}$. normal-$\square$
Non-trivial uncertainty sets using the Wasserstein distance or the $F$-norm, beyond the diagonal $\epsilon$-perturbation (cf. Example 6), for $\hat{\bm{R}}_{\underline{z}}$, $\hat{\bm{R}}_{\underline{xz}}$, and $\hat{\bm{R}}_{\underline{s}}$ can be straightforwardly employed and the distributional robustness modeling and analyses remain routine; see Subsection IV-B. Hence, we omit them here.

### v2

@@ IV-B2 F-Norm Distributionally Robust Beamforming
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{(\bm{R}-\hat{\bm{R}})^{\mathsf{H}}(\bm{R}-\hat{\bm{R}})}\right]\leq\epsilon^{2}_{0}\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0}.\end{array}$ (52)
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R},\bm{U}}&\operatorname{Tr}\bm{R}\\ \text{s.t.}&\operatorname{Tr}\left[{\bm{U}}\right]\leq\epsilon^{2}_{0},\\ &\left[\begin{array}[]{cc}\bm{U}&(\bm{R}-\hat{\bm{R}})^{\mathsf{H}}\\ (\bm{R}-\hat{\bm{R}})&\bm{I}_{N+M}\end{array}\right]\succeq\bm{0},\\ &\bm{R}\succeq\bm{0},~{}\bm{R}_{x}\succ\bm{0},~{}\bm{U}\succeq\bm{0}.\end{array}$ (53)
@@ IV-C Multi-Frame Case: Dynamic Channel Evolution
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{Wiener-MF}}&=[\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}}][\bm{R}_{x}+\lambda\bm{I}_{N}]^{-1}\\ &=[\bm{R}_{s}\bm{H}^{\mathsf{H}}+\lambda\bm{W}^{\prime\mathsf{H}}][\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}+\lambda\bm{I}_{N}]^{-1},\end{array}$ (55)
where $\lambda\geq 0$ is a tuning parameter to control the similarity between $\bm{W}$ and $\bm{W}^{\prime}$. Specifically, if $\lambda$ is large, $\bm{W}$ must be close to $\bm{W}^{\prime}$; if $\lambda$ is small, $\bm{W}$ can be far away from $\bm{W}^{\prime}$. $\square$
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\big{[}-(\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}})^{\mathsf{H}}\cdot(\bm{R}_{x}+\lambda\bm{I}_{N})^{-1}\cdot\\ &\quad\quad\quad(\bm{R}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}})+(\bm{R}_{s}+\lambda\bm{W}^{\prime}\bm{W}^{\prime\mathsf{H}})\big{]}\\ \text{s.t.}&d_{0}(\bm{R},{\hat{\bm{R}}})\leq\epsilon_{0},\\ &\bm{R}\succeq\bm{0},\end{array}$ (56)
whose objective function is monotonically increasing in $\bm{R}$.
The remaining distributional robustness modeling and analyses against the uncertainties in $\bm{R}$ are technically straightforward, and therefore, we omit them here. Upon using the diagonal-loading method on $\bm{R}$, a distributionally robust beamformer for the multi-frame case is
$\begin{array}[]{l}\bm{W}^{\star}_{\text{DR-Wiener-MF}}=[\hat{\bm{R}}_{xs}+\lambda\bm{W}^{\prime\mathsf{H}}]\cdot[\hat{\bm{R}}_{x}+\lambda\bm{I}_{N}+\epsilon_{0}\bm{I}_{N}]^{-1},\end{array}$
where $\epsilon_{0}$ is an uncertainty quantification parameter for $\bm{R}$.
@@ V-A1 General Framework and Concrete Examples
$\min_{\bm{W}\in\mathbb{R}^{2M\times L}}\max_{\mathbb{P}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}\in\mathcal{U}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}}\operatorname{Tr}\mathbb{E}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}[\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}][\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}]^{\mathsf{T}}.$ (57)
The proposition below reformulates and solves (57).
@@
Let $\underline{\mathbf{z}}\coloneqq\bm{\varphi}(\underline{\mathbf{x}})$. Then, the distributionally robust $\underline{\mathbf{x}}$-nonlinear estimation problem (57) can be rewritten as a distributionally robust $\underline{\mathbf{z}}$-linear beamforming problem as
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{\underline{z}}\bm{W}^{\mathsf{T}}-\bm{W}\bm{R}_{\underline{zs}}-\bm{R}^{\mathsf{T}}_{\underline{zs}}\bm{W}^{\mathsf{T}}+\bm{R}_{\underline{s}}\big{]}\\ \text{s.t.}&d_{0}\left(\left[\begin{array}[]{cc}\bm{R}_{\underline{z}}&\bm{R}_{\underline{zs}}\\ \bm{R}^{\mathsf{T}}_{\underline{zs}}&\bm{R}_{\underline{s}}\end{array}\right],\left[\begin{array}[]{cc}\hat{\bm{R}}_{\underline{z}}&\hat{\bm{R}}_{\underline{zs}}\\ \hat{\bm{R}}^{\mathsf{T}}_{\underline{zs}}&\hat{\bm{R}}_{\underline{s}}\end{array}\right]\right)\leq\epsilon_{0},\\ &\left[\begin{array}[]{cc}\bm{R}_{\underline{z}}&\bm{R}_{\underline{zs}}\\ \bm{R}^{\mathsf{T}}_{\underline{zs}}&\bm{R}_{\underline{s}}\end{array}\right]\succeq\bm{0},\end{array}$ (58)
where $\hat{\bm{R}}_{\underline{z}}=\frac{1}{L}\bm{K}^{2}$, $\hat{\bm{R}}_{\underline{zs}}=\frac{1}{L}\bm{K}\underline{\bm{S}}^{\mathsf{T}}$, $\hat{\bm{R}}_{\underline{s}}=\frac{1}{L}\underline{\bm{S}}\underline{\bm{S}}^{\mathsf{T}}$, and $\underline{\bm{S}}\coloneqq[\operatorname{Re}\bm{S};~{}\operatorname{Im}\bm{S}]=[\underline{\bm{s}}_{1},\underline{\bm{s}}_{2},\ldots,\underline{\bm{s}}_{L}]$. In addition, the strong min-max property holds for (58): i.e., the order of $\min$ and $\max$ can be exchanged provided that the first constraint is compact convex. As a result, given every pair of $(\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}})$, the optimal Wiener beamformer is [^3] For numerical stability in the inversion of $\bm{K}$, $L$ should not be too large.
$\bm{W}^{\star}_{\text{RKHS}}=\hat{\bm{R}}^{\mathsf{T}}_{\underline{zs}}\cdot\hat{\bm{R}}^{-1}_{\underline{z}}=\underline{\bm{S}}\bm{K}\bm{K}^{-2}=\underline{\bm{S}}\bm{K}^{-1},$ (59)
@@
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{\underline{z}},\bm{R}_{\underline{zs}},\bm{R}_{\underline{s}}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{T}}_{\underline{zs}}\bm{R}^{-1}_{\underline{z}}\bm{R}_{\underline{zs}}+\bm{R}_{\underline{s}}\big{]}\\ \text{s.t.}&d_{0}\left(\left[\begin{array}[]{cc}\bm{R}_{\underline{z}}&\bm{R}_{\underline{zs}}\\ \bm{R}^{\mathsf{T}}_{\underline{zs}}&\bm{R}_{\underline{s}}\end{array}\right],~{}\left[\begin{array}[]{cc}\hat{\bm{R}}_{\underline{z}}&\hat{\bm{R}}_{\underline{zs}}\\ \hat{\bm{R}}^{\mathsf{T}}_{\underline{zs}}&\hat{\bm{R}}_{\underline{s}}\end{array}\right]\right)\leq\epsilon_{0},\\ &\left[\begin{array}[]{cc}\bm{R}_{\underline{z}}&\bm{R}_{\underline{zs}}\\ \bm{R}^{\mathsf{T}}_{\underline{zs}}&\bm{R}_{\underline{s}}\end{array}\right]\succeq\bm{0},~{}~{}~{}\bm{R}_{\underline{z}}\succ\bm{0}.\end{array}$ (60)
@@
Treating $[\underline{\mathbf{z}};\underline{\mathbf{s}}]$ as, or approximating $[\underline{\mathbf{z}};\underline{\mathbf{s}}]$ using, a joint Gaussian random vector due to the linear estimation relation $\hat{\underline{\mathbf{s}}}=\bm{W}\underline{\mathbf{z}}$ in RKHS [cf. (57)], then the results in Lemma 1 apply. For details, see Appendix F of the online supplementary materials. $\square$
In (58), $d_{0}$ defines a matrix similarity measure to quantify the uncertainty of the covariance matrix of $[\underline{\mathbf{z}};\underline{\mathbf{s}}]$, and $\epsilon_{0}\geq 0$ quantifies the uncertainty level. Proposition 3 reveals the benefit of the kernel trick (2), that is, the possibility to represent a nonlinear estimation problem as a linear one.
@@
which upper bounds the true estimation error covariance. $\square$
@@
By using the trimmed diagonal-loading uncertainty set for $\bm{R}_{\underline{z}}$, i.e.,
$\hat{\bm{R}}_{\underline{z}}-\epsilon_{0}\bm{I}_{L}\preceq\bm{R}_{\underline{z}}\preceq\hat{\bm{R}}_{\underline{z}}+\epsilon_{0}\bm{I}_{L},$
@@
$\bm{\phi}^{\star}(\mathbf{x})=\bm{\Gamma}_{M}\cdot\frac{1}{L}\underline{\bm{S}}\bm{K}\cdot\left(\frac{1}{L}\bm{K}^{2}+\epsilon_{0}\bm{I}_{L}\right)^{-1}\cdot\bm{\varphi}(\underline{\mathbf{x}}),$ (63)
@@
$\begin{array}[]{l}\displaystyle\min_{\bm{W}}\operatorname{Tr}\mathbb{E}_{(\underline{\mathbf{x}},\underline{\mathbf{s}})\sim\hat{\mathbb{P}}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}}[\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}][\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}]^{\mathsf{T}}+\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\epsilon_{0}\cdot\operatorname{Tr}[\bm{W}\bm{W}^{\mathsf{T}}],\end{array}$ (64)
which can be proven by replacing $\bm{R}_{\underline{z}}$ in (58) with its upper bound. $\square$
@@
The kernelized eigenvalue thresholding method can be designed in analogy to Example 2. The two key steps are to obtain the eigenvalue decomposition of $\hat{\bm{R}}_{\underline{z}}=\bm{K}^{2}/L$ and then lift the eigenvalues; cf. (39). $\square$
@@
${\hat{\bm{R}}}_{\underline{z}}-\epsilon_{0}\bm{I}_{L}\preceq{\bm{R}}_{\underline{z}}\preceq{\hat{\bm{R}}}_{\underline{z}}+\epsilon_{0}\bm{I}_{L},$
@@
${\hat{\bm{R}}}_{\underline{z}}-\epsilon_{0}\bm{F}\preceq{\bm{R}}_{\underline{z}}\preceq{\hat{\bm{R}}}_{\underline{z}}+\epsilon_{0}\bm{F}$
for some $\bm{F}\succeq\bm{0}$.
@@
Theorem 3 gives the kernel ridge regression an interpretation of distributional robustness. The usual choice of $\bm{F}$ in Theorem 3 is the $L$-divided kernel matrix $\bm{K}/L$; see, e.g., [36, Eq. (4)], [24, Eqs. (15.110) and (15.113)]. As a result, from (63), we have
$\bm{\phi}^{\star}(\mathbf{x})=\bm{\Gamma}_{M}\cdot\underline{\bm{S}}\cdot\left(\bm{K}+\epsilon_{0}\bm{I}_{L}\right)^{-1}\cdot\bm{\varphi}(\underline{\mathbf{x}}),$ (65)
which is another type of kernel ridge regression (i.e., a new kernelized diagonal-loading method).
@@
$\begin{array}[]{l}\displaystyle\min_{\bm{W}}\operatorname{Tr}\mathbb{E}_{(\underline{\mathbf{x}},\underline{\mathbf{s}})\sim\hat{\mathbb{P}}_{\underline{\mathbf{x}},\underline{\mathbf{s}}}}[\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}][\bm{W}\cdot\bm{\varphi}(\underline{\mathbf{x}})-\underline{\mathbf{s}}]^{\mathsf{T}}+\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\lambda\cdot\|\bm{W}\|^{2},\end{array}$ (66)
@@
Consider the nonlinear regression problem in Theorem 3. Its data-perturbed counterpart can be constructed by taking into account the data perturbation vectors $(\mathbf{\Delta}_{\underline{s}},\mathbf{\Delta}_{\underline{z}})$. Suppose that $\mathbf{\Delta}_{\underline{z}}$ is uncorrelated with $\underline{\mathbf{z}}$, with $\underline{\mathbf{s}}$, and with $\mathbf{\Delta}_{\underline{s}}$; in addition, $\mathbf{\Delta}_{\underline{s}}$ is uncorrelated with $\underline{\mathbf{z}}$. If the second-order moment of $\mathbf{\Delta}_{\underline{z}}$ is upper bounded by $\epsilon_{0}\bm{I}_{L}$, then the distributionally robust estimator of $\bm{W}$ becomes a kernel ridge regression (i.e., squared-$F$-norm regularized) method (64). The regularization term becomes $\operatorname{Tr}\big{[}\bm{W}\bm{F}\bm{W}^{\mathsf{H}}\big{]}$, which is known as the Tikhonov regularizer, if the second-order moment of $\mathbf{\Delta}_{\underline{z}}$ is upper bounded by $\epsilon_{0}\bm{F}$ for some $\bm{F}\succeq\bm{0}$. $\square$
General uncertainty sets using the Wasserstein distance or the $F$-norm, beyond the diagonal $\epsilon_{0}$-perturbation (cf. Example 4), can be straightforwardly employed and the distributional robustness modeling and analyses remain routine; cf. Subsection IV-B. Hence, we omit them here. However, such complicated approaches are computationally prohibitive in practice when $L$ or $M$ is large.